help please and thank you
find the LCM of 7(v-2) and 14(v-2)

Found 2 solutions by NancyLam, richard1234:
Answer by NancyLam(35)

the LCM is: 7v-14
not sure

Answer by richard1234(7193)

LCM is the least common multiple of both numbers...the LCM must be an integer multiple of both numbers. Factoring,

7(v-2) = 7(v-2) in simplest form (we could expand as 7v - 14 but it's easier to assume v-2 is a factor)

14(v-2) = 2*7(v-2)

Taking the intersection of the factors, the LCM is 14(v-2).

(Intersection refers to the intersection of two sets, in this case, each factorization is a set).